# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v3

@@ IV-D Route Generation Module
Given predicted future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$, the Route Generation Module generates multiple possible future routes from these probabilities. An $n$-ary tree-based algorithm, Spanning Route, is proposed to generate these future routes based on the predicted probabilities. This algorithm is visualized in Figure 5 through a simplified case where $n=2$ and only $\Gamma=3$ predicted future links are illustrated. For each tree node, we designate four attributes: name, parent, end_node, and pred. The name corresponds to the identification of the leaf, while the parent points to the predecessor of the current leaf. The attribute end_node signifies the terminal node of the current predicted link, and pred is the present predictions.
@@
3root $\leftarrow$ CreateNewNode(name = “root”, parent = NIL, end_node = $v_{i,\Gamma}^{s}$, pred = NIL)
@@
9             $\left\{e_{i,k}^{\Gamma+\gamma}\right\}_{k=1}^{n}$ = GetTopK($\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})[\mathcal{N}_{end\_node}^{e}]$, $K=n$)
@@
11                   node = CreateNewNode(name = “$k$”, parent = leaf, end_node = $e_{i,k}^{\Gamma+\gamma}[1]$, pred = $e_{i,k}^{\Gamma+\gamma}$)
@@ 16       $\text{path}_{k}$ = Traverse(root, leaves[$k$])
To formally introduce the Spanning Route algorithm, we provide the pseudo-code for generating the multiple future routes based on the predicted probabilities in Algorithm 1. Specifically, the algorithm encompasses four primary functions. The CreateNewNode function instantiates a new node in the tree given its attributes, while the GetLeaves function takes the root node as input and outputs all leaves of the tree. The GetTopK function retrieves the top-$k$ predictions given a predicted probability and $k$, and the Traverse function applies a tree-based Depth-First Search (DFS) traversal algorithm, specifically a Pre-Order traversal [76]—to acquire the path from the root to a specified leaf. This last function is instrumental in merging the predicted links into cohesive predicted routes. We note that certain detailed masking and indexing operations have been omitted in the presented pseudo code for clarity. For more details, please refer to the minibatch version of the Spanning Route algorithm 3 detailed in the appendix (VI-A).
@@ IV-E Rank Refinement Module
The top-$K$ future routes candidates $\widetilde{\mathcal{X}^{f}}=\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\{\{\widetilde{x_{i,k}^{f}}\}_{k=1}^{K}\}_{i=1}^{|\mathcal{X}^{f}|}$ offer an initial selection of possible future routes. However, the dependencies among different links within these routes are solely based on the connectivity of the road network. Given that the consistency and other spatial relations of links within a route also affect people’s choices of routes, a more refined approach is needed for accurate future route prediction. To achieve this, we leverage learned spatial relations $\mathcal{R}$ to model the dependencies between different links and rerank the candidate routes based on the learned dependencies. This process can be denoted as $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$
Consider the future routes $\widetilde{x_{i}^{f}}\in\mathbb{R}^{K\times\Gamma^{\prime}}$. Initially, these routes are encoded by leveraging the embedding matrices $\mathbf{W}_{\mathcal{E}}$ and $\mathbf{W}_{\mathcal{R}^{d}}$, thereby resulting in the route embedding $\widetilde{\mathbf{x}^{f}_{i}}\in\mathbb{R}^{K\times\Gamma^{\prime}\times\delta_{\mathcal{E}}}$ and route direction embedding $\widetilde{\mathbf{x}^{f,d}_{i}}\in\mathbb{R}^{K\times\Gamma^{\prime}\times\delta_{\mathcal{R}^{d}}}$. In the subsequent reranking phase, we prioritize the routes with higher consistency and connectivity, utilizing the spatial relations $\mathcal{R}$ learned from the Knowledge Graph Module. Specifically, the obtained route embeddings are projected onto the “ConnectBy” and “ConsistentWith” hyperplanes as follows:
@@
where $\mathbf{r}^{f,c}_{i,m}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ is the connectivity margin and $\mathbf{r}^{f,s}_{i,m}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ the consistency margin. Following this, the derived $\mathbf{r}^{f,c}_{m}$ and $\mathbf{r}^{f,s}_{m}$ are flattened and, together with the flattened route embedding $\widetilde{\mathbf{x}^{f}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$ and route direction embedding $\widetilde{\mathbf{x}^{f,d}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$, used to compute the new rank:
$\mathrm{Pr}(\widetilde{R})=\mathrm{Softmax}(\mathrm{MLP}_{r}(\mathbf{r}^{c}_{m}\text{ }\|\text{ }\mathbf{r}^{s}_{m}\text{ }\|\text{ }\mathrm{MLP}_{f}(\widetilde{\mathbf{x}^{f}_{i}}\text{ }\|\text{ }\widetilde{\mathbf{x}^{f,d}_{i}}))),$ (13)
@@
$\widetilde{x^{f}_{i}}=\mathrm{MLP}_{k}(\mathbf{e}^{\Gamma}_{i}\text{ }\|\text{}\mathbf{e}^{d,\Gamma}_{i}\text{ }\|\text{ }\mathbf{r}^{d}_{i}\text{ }\|\text{ }\mathrm{MLP}_{x}(\mathbf{x}^{f}_{i}\text{ }\|\text{ }\mathbf{x}^{f,d}_{i})),$ (15)
@@ IV-F Multi-Objectives Optimization
The objective of RouteKG is to leverage the learned spatial relations to efficiently make future route predictions, which is done by optimizing multiple objectives. Specifically, the overall loss function could be written as:
@@
where $w_{\cdot}$ are weights for different loss. The complete learning procedure of RouteKG are detailed in Algorithm 2.
@@
Road network $\mathbf{G}$;
@@
NAE matrix $\mathbf{A}$;
1 Randomly initialize $\Theta_{kg}=\{\mathbf{W}_{\mathcal{E}},\mathrm{MLP}_{d},$ $\mathbf{W}_{\mathcal{R}^{c/s/a/d}},$$\mathbf{P}_{\mathcal{R}^{c/s/a/d}}\}\quad\text{and}\quad\Theta_{r}=\{\mathrm{MLP}_{r/f/k/x}\}.$
2 for $m=1,\dots,max\_iterations$ do
3       // Early stopping here.
4       Normalize embeddings of hyperplanes $\left\|\mathbf{P}_{\mathcal{R}^{\cdot}}\right\|_{2}=1$.
5       // Forward Propagation.
6       $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R}\leftarrow\mathcal{M}_{kg}\left(\mathcal{X}^{o},\mathbf{G},\mathbf{D};\Theta_{kg}\right)$
7       $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}\leftarrow\mathcal{M}_{g}(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A})$
8       $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}\leftarrow\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$
9       // Back Propagation.
10       $\Theta_{kg}\leftarrow\Theta_{kg}-\nabla_{\Theta_{kg}}\left\{\mathcal{L}_{rep}+\mathcal{L}_{d}+\mathcal{L}_{pred}\right\}$
11       $\Theta_{r}\leftarrow\Theta_{r}-\nabla_{\Theta_{r}}\left\{\mathcal{L}_{rank}\right\}$
12
Algorithm 2 RouteKG algorithm.
@@ V-A Data
Nodes Edges MID / MOD Max ID/OD Min ID/OD Density
Chengdu 2832 6506 2.297 4 0 8.11e-4
Shanghai 320 714 2.231 4 1 6.99e-3
TABLE II: Summary statistics of Chengdu and Shanghai road networks (ID: in-degree, OD: out-degree).
@@ V-B Baseline Methods
In this study, we compare our approach with several established baselines to evaluate performance:
  • Markov: The Markov model is a well-known sequential prediction method extensively used in the field. It bases its route forecasting on observed transition patterns between road links.
  • Dijkstra [37]: Dijkstra’s algorithm is a prominent method for finding the shortest paths between nodes in a graph, where the path length is assumed to be the sum of link length to reflect realistic geographic distance. This baseline can only work when the exact goal location is given.
@@
  • GRU [10]: The GRU is a type of RNN that utilizes gating mechanisms to capture long-term dependencies in the data, thereby improving the model’s predictive capabilities. This is especially beneficial for applications in trajectory prediction, where long-term dependencies play a crucial role.
  • LSTM [9]: The LSTM is another variant of RNNs that addresses the problem of learning long-term dependencies in data. It does this by implementing a special architecture consisting of a series of memory cells, which effectively control the flow of information, making LSTMs advantageous for predicting future routes.
  • NetTraj [14]: NetTraj is an advanced network-based trajectory prediction model specifically designed for predicting future movements in road networks. It leverages the inherent structure of road networks and utilizes historical trajectory data for accurate predictions. By integrating the Graph Attention Network (GAT) with LSTM, NetTraj offers a robust framework for making future trajectory predictions.
  • RCM-BC [77]: The Route Choice Model-Behavioral Cloning (RCM-BC) is a behavioral cloning approach designed for route choice modeling in sequential decision-making scenarios. It employs supervised learning to create a policy that maps states to actions based on observed behavior to predict future routes. This baseline also requires the knowledge of the exact goal location to work.
@@ V-C1 Experimental Settings
To comprehensively assess model performance, we design experiments based on the three subproblems as defined in Section III-A: (1) route prediction with unknown goal $\mathcal{F}_{1}$, (2) route prediction with goal direction only $\mathcal{F}_{2}$, and (3) route prediction with complete goal information $\mathcal{F}_{3}$. We refer to these three subproblems as NoGoal, GoalD, and Goal. They reflect varying degrees of information availability regarding the road user’s intended destination, and represent a broad range of real-world application scenarios. For instance, a system might not know the user’s exact destination due to privacy concerns but could have access to more general information, such as the goal direction.
@@
Most of the baseline models are designed for the NoGoal scenario, but two of them (Dikstra and RCM-BC) are for the Goal scenario only. Unlike these baselines, RouteKG requires the goal direction information. Therefore, specific model implementations are needed to incorporate the available goal information into different models under different scenarios. Under the NoGoal scenario, the goal direction is unknown, but we can still estimate it based on the observed route. Consequently, the estimated goal direction is used in RouteKG under NoGoal. Under GoalD, the actual goal direction is used instead of the estimated one in RouteKG. For other deep learning baseline models (except for RCM-BC), we concatenate the embedding of goal directions with the respective model’s inputs. Similarly, under the Goal scenario, the same concatenation strategy can be used for the baseline models, enriching them with complete goal information. In RouteKG, we add the embedding of the goal location directly to the embedding of the last link in the observed route.
@@
In our main experiments, the input observed route length is set as $\Gamma=10$ and output future route length as $\Gamma^{\prime}=5$. For model evaluation, the datasets are partitioned into training, validation, and test subsets in a 6:2:2 ratio. Different models are evaluated under the NoGoal, GoalD, and Goal scenarios, using both the “link-level” and “route-level” metrics. Link-level assessment has practical implications, particularly for tasks related to traffic flows, while route-level evaluation offers valuable information for routing applications. Specifically, we utilize Recall and Mean Reciprocal Rank (MRR), two prevalent metrics. Recall measures the ratio of relevant items retrieved from all relevant items, indicating the system’s capacity to fetch desired information. MRR, on the other hand, evaluates the rank position of the correct answer, computing the average reciprocal rank of the highest-ranked correct answer across queries. A higher MRR signifies superior performance. These metrics provide insights into model effectiveness and ranking quality and are useful tools for assessing and enhancing system performance.
@@
To define evaluation metrics, we consider the top-$k$ predictions for the $i$-th observed route $\{\widehat{x^{f}_{i,k}}\}_{k=1}^{K}$=$\{\{\widehat{e_{i,k}^{j}}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\}_{k=1}^{K}$ and the actual $i$-th future route $x_{i}^{f}=\{e_{i}^{j}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$. The link-level recall R@K is defined as $R@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\max_{k=1}^{K}\left[\frac{1}{\Gamma^{\prime}}\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}(\widehat{e_{i,k}^{j}}=e_{i}^{j})\right].$ where $\mathbb{I}(a=b)=\begin{cases}1&\text{if }a=b\\ 0&\text{otherwise}\end{cases}$ is the indicator function.
@@
The experiments are conducted on a Ubuntu server with the Python 3.6 environment. The deep learning computations are performed using the PyTorch framework. The server’s hardware specifications include an Intel(R) Xeon(R) Platinum 8375C CPU with a clock speed of 2.90GHz, coupled with 8 NVIDIA GeForce RTX 3090 GPUs, each featuring 24GB of memory. To ensure the robustness and generalizability of our model, hyperparameters are tuned based on the performance of the validation set. The fine-tuning of these hyperparameters is crucial for balancing the bias-variance trade-off and optimizing the model’s performance. All hyperparameters are in VI-B.

### v4

@@ IV-D Route Generation Module
[…]
@@
31ex
4root $\leftarrow$ CreateNewNode(name = “root”, parent = NIL, end_node = $v_{i,\Gamma}^{s}$, pred = NIL)
@@
[…]
@@
11         node = CreateNewNode(name = “kk”, parent = leaf, end_node = $e_{i,k}^{\Gamma+\gamma}[1]$, pred = $e_{i,k}^{\Gamma+\gamma}$)
@@ 16 $\text{path}_{k}$ = Traverse(root, leaves[$k$])
To formally introduce the Spanning Route algorithm, we provide the pseudo-code for generating multiple future routes based on the predicted probabilities in Algorithm 1. Specifically, the algorithm encompasses four primary functions. The CreateNewNode function instantiates a new node in the tree given its attributes, while the GetLeaves function takes the root node as input and outputs all leaves of the tree. The GetTopK function retrieves the top-$k$ predictions given a predicted probability and $k$, and the Traverse function applies a tree-based Depth-First Search (DFS) traversal algorithm, specifically a Pre-Order traversal [74]—to acquire the path from the root to a specified leaf. This last function is instrumental in merging the predicted links into cohesive predicted routes. We note that certain detailed masking and indexing operations have been omitted in the presented pseudo-code for clarity. For more details, please refer to the minibatch version of the Spanning Route algorithm 3 in the Appendix VI-A.
@@ IV-E Rank Refinement Module
The top-$K$ future routes candidates $\widetilde{\mathcal{X}^{f}}=\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\{\{\widetilde{x_{i,k}^{f}}\}_{k=1}^{K}\}_{i=1}^{|\mathcal{X}^{f}|}$ offer an initial selection of possible future routes. However, the dependencies among different links within these routes are solely based on the connectivity of the road network. Given that the consistency and other spatial relations of links within a route also affect people’s choices of routes, a more refined approach is needed for accurate future route prediction. To achieve this, we leverage learned spatial relations $\mathcal{R}$ to model the dependencies between different links and rerank the candidate routes based on the learned dependencies. This process can be denoted as $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$ .
Consider the future routes $\widetilde{x_{i}^{f}}\in\mathbb{R}^{K\times\Gamma^{\prime}}$. Initially, these routes are encoded using embedding matrices $\mathbf{W}_{\mathcal{E}}$ and $\mathbf{W}_{\mathcal{R}^{d}}$, thereby resulting in the route embedding $\widetilde{\mathbf{x}^{f}_{i}}\in\mathbb{R}^{K\times\Gamma^{\prime}\times\delta_{\mathcal{E}}}$ and route direction embedding $\widetilde{\mathbf{x}^{f,d}_{i}}\in\mathbb{R}^{K\times\Gamma^{\prime}\times\delta_{\mathcal{R}^{d}}}$. In the subsequent reranking phase, we prioritize the routes with higher consistency and connectivity, utilizing the spatial relations $\mathcal{R}$ learned from the Knowledge Graph Module. Specifically, the obtained route embeddings are projected onto the “ConnectBy” and “ConsistentWith” hyperplanes as follows:
@@
where $\mathbf{r}^{f,c}_{i,m}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ is the connectivity margin and $\mathbf{r}^{f,s}_{i,m}\in\mathbb{R}^{K\times\delta_{\mathcal{E}}}$ the consistency margin. These equations project candidate routes onto relation-specific hyperplanes (e.g., for ConnectBy and ConsistentWith), allowing the model to assess how well each route follows the learned spatial relations. Routes more aligned with these hyperplanes are considered more plausible and prioritized during reranking. Following this, the derived $\mathbf{r}^{f,c}_{m}$ and $\mathbf{r}^{f,s}_{m}$ are flattened and, together with the flattened route embedding $\widetilde{\mathbf{x}^{f}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$ and route direction embedding $\widetilde{\mathbf{x}^{f,d}_{i}}\in\mathbb{R}^{K\cdot\Gamma^{\prime}\cdot\delta_{\mathcal{E}}}$, used to compute the new rank:
[…]
@@
[…]
@@ IV-F Multi-Objectives Optimization
The objective of RouteKG is to leverage learned spatial relations to predict future routes. This is achieved through a multi-objective learning framework, combining several complementary goals: (1) $\mathcal{L}_{rep}$ encourages high-quality KG embeddings by modeling multiple spatial relations; (2) $\mathcal{L}_{d}$ ensures accurate direction classification for the EGoalD setting; (3) $\mathcal{L}_{pred}$ guides the prediction of future links via KGC; (4) $\mathcal{L}_{rank}$ improves final route quality by reranking based on relational consistency. Collectively, these objectives support each stage of the prediction pipeline, ensuring that the model benefits from both structural knowledge and task-specific supervision. These objectives are jointly optimized through the following weighted total loss:
@@
where $w_{\cdot}$ are the balancing weights. The overall training procedure that integrates all components is summarized in Algorithm 2.
1
@@
Road network $\mathbf{G}$; NAE matrix $\mathbf{A}$;
@@
2
31exInit: Randomly initialize KG parameters $\Theta_{kg}$ and rerank parameters $\Theta_{r}$.;
4 for $m=1,\dots,\textit{max\_iters}$ do
5 // — Step 1: KG representation learning —
6   Normalize hyperplane embeddings $\left\|\mathbf{P}_{\mathcal{R}^{\cdot}}\right\|_{2}=1$;
7   Sample positive/negative triplets $\Delta,\Delta^{\prime}$ from $\mathcal{X}^{o}_{\mathcal{B}}$ and $\mathbf{G}$;
8   Compute $\mathcal{L}_{rep}$ (Eq. 4);
9
10  1ex
11 // — Step 2: forward prediction —
12 if setting = NoGoal then
13    $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R},\widehat{r^{d}}\leftarrow\mathcal{M}_{kg}\!\bigl(x^{o},\,\mathbf{G},\mathbf{D};\Theta_{kg}\bigr)$
14      Compute $\mathcal{L}_{d}$ (Eq. 6)
15 else if setting = GoalD then
16    $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R}\leftarrow\mathcal{M}_{kg}\!\bigl(\{x^{o},r^{d}\},\mathbf{G},\mathbf{D};\Theta_{kg}\bigr)$
17 else if setting = Goal then
18    $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R}\leftarrow\mathcal{M}_{kg}\!\bigl(\{x^{o},r^{d},e^{\Gamma+\Gamma^{\prime}}\},\mathbf{G},\mathbf{D};\Theta_{kg}\bigr)$
19
20  1ex// — Step 3: spanning-route + rerank —
21   Candidate routes $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}\leftarrow\mathcal{M}_{g}\bigl(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A}\bigr)$;
22   Compute $\mathcal{L}_{pred}$ (Eq. 10);
23   Ranked routes $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}\leftarrow\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\},\mathcal{R};\Theta_{r})$;
24   Compute $\mathcal{L}_{rank}$ (Eq. 14)
25  1ex// — Step 4: parameter update —
26 $\Theta_{kg}\!\leftarrow\!\Theta_{kg}-\eta\nabla_{\Theta_{kg}}(\mathcal{L}_{rep}+\mathcal{L}_{d}+\mathcal{L}_{pred})$;
27 $\Theta_{r}\!\leftarrow\!\Theta_{r}-\eta\nabla_{\Theta_{r}}(\mathcal{L}_{rank})$;
28
Algorithm 2 RouteKG training procedure.
@@ V-A Data
TABLE II: Statistics of Chengdu and Shanghai road networks. (M)ID: (Mean) in-degree, (M)OD: (Mean) out-degree.
Nodes Edges MID / MOD Max ID/OD, Min ID/OD Density
Chengdu 2832 6506 2.297 4, 0 8.11e-4
Shanghai 320 714 2.231 4, 1 6.99e-3
@@ V-B Baseline Methods
In this study, we compare our approach with several established baselines to evaluate model performance:
  • Markov: The Markov model is a widely-used sequential prediction method. It bases its route forecasting on observed transition patterns between road links.
  • Dijkstra [35]: Dijkstra’s algorithm is a prominent method for finding the shortest path in a graph, where the path length is assumed to be the sum of link lengths. This baseline can only work when the exact goal location is given.
@@
  • GRU [9]: The GRU is a type of RNN that utilizes gating mechanisms to capture long-term dependencies in the data, thereby improving the model performance for trajectory or route prediction.
  • LSTM [8]: As another variant of RNNs, the LSTM learns long-term dependencies in data by implementing a special architecture consisting of a series of memory cells, which effectively control the flow of information.
  • NetTraj [13]: NetTraj is an advanced network-based trajectory prediction model specifically designed for predicting future movements in road networks. By integrating the Graph Attention Network (GAT) with LSTM, it leverages the spatial structure of road networks and historical trajectory data for robust future trajectory predictions.
  • RCM-BC [75]: RCM-BC is a behavioral cloning approach designed for route choice modeling in sequential decision-making scenarios. It employs supervised learning to create a policy that maps states to actions based on observed behavior to predict future routes. This baseline also requires knowledge of the exact goal location.
  • RoPT [76]: RoPT combines stacked GCN layers with a Transformer encoder to capture spatial and long-range temporal dependencies.
TABLE III: Performance comparison of different methods.
@@ V-C1 Experimental Settings
To comprehensively assess model performance, we design experiments based on the three subproblems as defined in Section III-A: (1) route prediction with unknown goal $\mathcal{F}_{1}$, (2) route prediction with goal direction only $\mathcal{F}_{2}$, and (3) route prediction with complete goal information $\mathcal{F}_{3}$. We refer to these three subproblems as NoGoal, GoalD, and Goal. They reflect varying degrees of information availability regarding the road user’s intended destination, and represent a broad range of real-world application scenarios. For instance, a system might not know the user’s exact destination due to privacy concerns, but could have access to more general information, such as the goal direction.
@@
Most of the baseline models are designed for the NoGoal scenario, but two of them (Dijkstra and RCM-BC) are for the Goal scenario only. Unlike these baselines, RouteKG requires the goal direction information. Therefore, specific model implementations are needed to incorporate the available goal information into different models under different scenarios. Under the NoGoal scenario, the goal direction is unknown, but we can still estimate it based on the observed route. Consequently, the estimated goal direction is used in RouteKG under NoGoal. Under GoalD, the actual goal direction is used instead of the estimated one in RouteKG. For other deep learning baseline models (except for RCM-BC), we concatenate the embedding of goal directions with the respective model’s inputs. Similarly, under the Goal scenario, the same concatenation strategy can be used for the baseline models, enriching them with complete goal information. In RouteKG, we add the embedding of the goal location directly to the embedding of the last link in the observed route. The ConsistentWith relation in the KG is constructed using observed trajectories to compute link co-occurrence. This requires access to full observed trajectories $\mathcal{X}^{o}$ during training, but not at inference time. For the EGoalD scenario, both training and inference require the observed route prefix $\mathcal{X}^{o}$ to estimate the goal direction. In contrast, under the GoalD and Goal settings, both training and inference rely only on the last observed link $e^{\Gamma}$ to predict the future route. This makes RouteKG applicable even in cases with limited route history.
@@
In our main experiments, the input observed route length is set as $\Gamma=10$ and the output future route length as $\Gamma^{\prime}=5$. For model evaluation, the datasets are partitioned into training, validation, and test subsets in a 6:2:2 ratio. Different models are evaluated under the NoGoal, GoalD, and Goal scenarios, using both the “link-level” and “route-level” metrics. Link-level assessment has practical implications, particularly for tasks related to traffic flows, while route-level evaluation offers valuable information for routing applications. Specifically, we utilize Recall and Mean Reciprocal Rank (MRR), two prevalent metrics. Recall measures the ratio of relevant items retrieved from all relevant items, indicating the system’s capacity to fetch desired information. MRR, on the other hand, evaluates the rank position of the correct answer, computing the average reciprocal rank of the highest-ranked correct answer across queries. A higher MRR signifies superior performance. These metrics provide insights into model effectiveness and ranking quality and are useful tools for assessing and enhancing system performance.
@@
To define evaluation metrics, we consider the top-$k$ predictions for the $i$-th observed route $\{\widehat{x^{f}_{i,k}}\}_{k=1}^{K}$$\{\{\widehat{e_{i,k}^{j}}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\}_{k=1}^{K}$ and the actual $i$-th future route $x_{i}^{f}=\{e_{i}^{j}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$. The link-level recall R@K is defined as $R@K=\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\max_{k=1}^{K}\left[\frac{1}{\Gamma^{\prime}}\sum_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}\mathbb{I}(\widehat{e_{i,k}^{j}}=e_{i}^{j})\right].$ where $\mathbb{I}(a=b)=\begin{cases}1&\text{if }a=b\\ 0&\text{otherwise}\end{cases}$ is the indicator function.
@@
The experiments are conducted on a Ubuntu server with the Python 3.6 environment. The deep learning computations are performed using the PyTorch framework. The server’s hardware specifications include an Intel(R) Xeon(R) Platinum 8375C CPU with a clock speed of 2.90GHz, coupled with 8 NVIDIA GeForce RTX 3090 GPUs, each featuring 24GB of memory. To ensure the robustness and generalizability of our model, hyperparameters are tuned based on the performance of the validation set, which is crucial for balancing the bias-variance trade-off and optimizing the model’s performance. All hyperparameter settings are detailed in Appendix VI-B.